# Non-Hermitian Linear Electro-Optic Effect Through Interactions of Free and Bound Charges

### v1

@@ C.1 Solutions of Maxwell equations
×$\nabla\times\mathbf{E}$ $=i\omega\mu_{0}\mathbf{H}$ (52)
×$\nabla\times\mathbf{H}$ $=-i\omega\varepsilon_{0}\mathbf{\overline{\varepsilon}}\cdot\mathbf{E}$ (53)
@@
where the $4\times 4$ matrix $\text{e}^{-iy\mathbf{\overline{M}}}$ is obtained by taking the exponential of the matrix of $-iy\mathbf{\overline{M}}$. Equation (56) relates the state vectors $\mathbf{f}(y)$ and $\mathbf{f}(0)$ at two distinct points of the material in terms of the transfer matrix $\text{e}^{-iy\mathbf{\overline{M}}}$.
@@ C.2 Reflection and transmission matrices
$\mathbf{H}_{t}^{\pm}$ =±$=\pm\frac{1}{\eta_{0}}\mathbf{J}\cdot\mathbf{Y}_{\text{ext}}\cdot\mathbf{E}_{t}^{\pm},$ (57)
@@
with $\eta_{0}$ the vacuum impedance and $k_{y}=\sqrt{\dfrac{\omega^{2}\varepsilon_{\text{ext}}}{c^{2}}-k_{x}^{2}-k_{z}^{2}}$. In Eq. (57) the superscript $\pm$ discriminates between plane waves propagating along the $+$ or $-y$ direction, respectively.

### v2

@@ C.1 Solutions of Maxwell equations
$\nabla\times\mathbf{E}$ $=i\omega\mu_{0}\mathbf{H}$ (52)
$\nabla\times\mathbf{H}$ $=-i\omega\varepsilon_{0}\mathbf{\overline{\varepsilon}}\cdot\mathbf{E}$ (53)
@@
where the $4\times 4$ matrix $\text{e}^{-iy\mathbf{\overline{M}}}$ is obtained by taking the exponential of the matrix of $iy\mathbf{\overline{M}}$. Equation (56) relates the state vectors $\mathbf{f}(y)$ and $\mathbf{f}(0)$ at two distinct points of the material in terms of the transfer matrix $\text{e}^{-iy\mathbf{\overline{M}}}$.
@@ C.2 Reflection and transmission matrices
$\mathbf{H}_{t}^{\pm}$ $=\pm\frac{1}{\eta_{0}}\mathbf{J}\cdot\mathbf{Y}_{\text{ext}}\cdot\mathbf{E}_{t}^{\pm},$ (57)
@@
with $\eta_{0}$ the vacuum impedance and $k_{y}=\sqrt{\dfrac{\omega^{2}\varepsilon_{\text{ext}}}{c^{2}}-k_{x}^{2}-k_{z}^{2}}$. In Eq. (57) the superscript $\pm$ discriminates between plane waves propagating along the ++ or y-y direction, respectively.